OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a red marble?

OpenStudy (kropot72):

On the first selection the probability of green is \[\frac{4}{10}\] On the second selection, having picked green on the first, the probability of red is \[\frac{3}{9}\] The probability of green then red is \[\frac{4}{10} \times\frac{3}{9}=you\ can\ calculate\]

OpenStudy (anonymous):

I got 2/15

OpenStudy (anonymous):

Thanks @kropot72

OpenStudy (kropot72):

Yes @WAyne5 your answer is correct :) You're welcome!